Simplify the radical expressions. Assume that all variables represent positive real numbers.
/25 + /48 + /49 + /192

Answer by checkley71(8403)

sqrt25=5
sqrt48=4sqrt3
sqrt49=7
sqrt192=8sqrt3
we now have:
5+7+4sqrt3+8sqrt3=12+12sqrt3=12(1+sqrt3) answer.